OpenStudy (anonymous):

2x^2+2x=4 I got to two separate answers, 2 & √5 -1 Which one is truly the answer?

OpenStudy (anonymous):

plug them back into the original equation: 2x^2+2x and see which one gives you back 4.

OpenStudy (anonymous):

Neither of these..

OpenStudy (anonymous):

2x^2+2x=4 divide both sides by 2 x^2+x=4 x^2+x-4 = 0 solve using the quadratic formula

OpenStudy (anonymous):

2x^2 + 2x = 4 2x^2 + 2x -4 = 0 2 ( x^2 + x - 2) = 0 2 (x-1) (x+2) = 0 x-1=0 : x=1 x+2=0 : x=-2

OpenStudy (anonymous):

I see what I did wrong! Thanks so much!

OpenStudy (comm.dan):

@Virgo91684 It is just basically Factoring by Grouping for polynomials

OpenStudy (comm.dan):

 @Virgo91684

OpenStudy (anonymous):

Oh the infinity symbol

OpenStudy (comm.dan):

Thank you for that reminder @Virgo91684

OpenStudy (comm.dan):

Always remember this for when finding domain from something, Domain is always all real numbers @Virgo91684

OpenStudy (anonymous):

So no negatives?

OpenStudy (comm.dan):

Nope not ever, that is what I remember my teacher said to me. @Virgo91684

OpenStudy (comm.dan):

The first answer for the range that you sent me might be wrong, I just need the original problem please. @Virgo91684

OpenStudy (anonymous):

f(x) = -sqrt(x-3)

OpenStudy (comm.dan):

What does the -sqrt stand for? @Virgo91684

OpenStudy (anonymous):